# Optimal Reward Labeling: Bridging Offline Preference and Reward-Based Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning has become one of the most practical RL settings. A recent success story has been RLHF, offline preference-based RL (PBRL) with preference from humans. However, most existing works on offline RL focus on the standard setting with scalar reward feedback. It remains unknown how to universally transfer the existing rich understanding of offline RL from the reward-based to the preference-based setting. In this work, we propose a general framework to bridge this gap. Our key insight is transforming preference feedback to scalar rewards via optimal reward labeling (ORL), and then any reward-based offline RL algorithms can be applied to the dataset with the reward labels. We theoretically show the connection between several recent PBRL techniques and our framework combined with specific offline RL algorithms in terms of how they utilize the preference signals. By combining reward labeling with different algorithms, our framework can lead to new and potentially more efficient offline PBRL algorithms. We empirically test our framework on preference datasets based on the standard D4RL benchmark. When combined with a variety of efficient reward-based offline RL algorithms, the learning result achieved under our framework is comparable to training the same algorithm on the dataset with actual rewards in many cases and better than the recent PBRL baselines in most cases.  [^1] https://github.com/uiuc-focal-lab/ORL
@@ 1 Introduction
Reinforcement learning (RL) is an important learning paradigm for solving sequential decision-making problems (Sutton and Barto, 2018). Compared to the standard (reward-based) RL that requires access to reward feedback (Schulman et al., 2017), preference-based RL (PBRL) (Wirth et al., 2017) only requires preference feedback over a pair of trajectories, making it more accessible in practice. In the offline learning setting, the agent only needs a pre-collected dataset of preference labels before training, making it even more convenient (Zhu et al., 2023). When humans provide preference labels, offline PBRL is known as reinforcement learning from human feedback (RLHF) (Stiennon et al., 2020), a popular framework for aligning large language models.
Despite the success achieved by offline PBRL (Kim et al., 2023; Dai et al., 2023; Hejna and Sadigh, 2024), the topic has been studied less than standard offline RL (Li et al., 2024; Levine et al., 2020; Fujimoto and Gu, 2021; Kidambi et al., 2020; Wu et al., 2019; Cheng et al., 2022; Tarasov et al., 2024). The critical feature in offline learning is ‘pessimism’ (Li et al., 2024). The learning agent should be pessimistic about the policies whose behaviors are not included in the dataset. Many pessimistic learning algorithms with theoretical insights have been developed for the standard offline RL setting. However, most empirical works for offline PBRL only adopt limited specific pessimistic learning approaches, such as applying certain regularization to the learning policy (Rafailov et al., 2024; Stiennon et al., 2020), which are empirically shown to be less efficient compared to recent SOTA methods in the standard RL setting (Cheng et al., 2022; Tarasov et al., 2024).
Figure 1: Illustration of the optimal reward labeling (ORL) framework. Given a dataset consisting of preference feedback as input, the ORL framework labels the dataset with the rewards that best explain the preference signals from the dataset. With a dataset of reward labels, the learning agent can train on the dataset with any efficient offline RL algorithms such as IQL (Kostrikov et al., 2021) or CQL (Kumar et al., 2020) and enjoy high learning efficiency.
@@
  • Standard offline RL algorithms are very different from each other. Model-based algorithms (Kidambi et al., 2020; Yu et al., 2020) learn an environment model and then run online learning algorithms on the model. Model-free algorithms (Cheng et al., 2022; Fujimoto and Gu, 2021) learn a pair of actor-critic models directly to infer the optimal policy without modeling the environment. The framework needs to apply to all kinds of learning algorithms.
  • Preference feedback contains arguably less information than reward feedback. One cannot recover the actual rewards from the preference feedback. The framework needs to make the standard offline RL algorithms work with incomplete information from preference feedback.
In this work, we build a novel framework called ‘ORL’ based on a reward labeling technique that labels the preference dataset with scalar rewards. The critical insight is that we only need to consider transforming feedback signals from preference to reward format while minimizing information loss. The standard offline RL algorithm can handle the pessimistic learning afterward. Note that reward labeling fundamentally differs from reward modeling, which is widely considered in recent PBRL studies (Kim et al., 2023; Stiennon et al., 2020). Reward modeling aims to infer rewards at state actions that are not in the preference dataset. In contrast, reward labeling focuses on interpreting the rewards for the state actions inside the preference dataset. To minimize the information loss in the transformation, we use the optimal reward labels that best explain the state actions in the dataset. We show that finding the optimal reward labels requires solving some simple optimization problems without needing complex function approximation. In Fig 1, we illustrate how our framework is used. In summary, our contributions are as follows.
  • We propose a general framework to bridge the gap between offline PBRL and standard offline RL. Our framework involves labeling the dataset with the optimal reward that best explains the preference dataset. Afterward, one can train on the reward-labeled dataset with any standard offline RL algorithms. Our framework can be easily implemented without modifying the original RL algorithms. One can possibly construct more efficient PBRL algorithms by applying SOTA standard RL algorithms to our framework.
@@ 2.1 Offline Reward-Based Reinforcement Learning
Offline reward-based reinforcement learning, or standard offline RL, is the most popular offline RL setting. The fundamental challenge for offline RL is known as ‘distribution mismatch’ (Levine et al., 2020). It refers to the phenomenon that the data distribution in the dataset may not match the distribution induced by the optimal policy. Due to distribution mismatch, an efficient learner must be ‘pessimistic’ (Li et al., 2024). In general, it says that an agent should rely more on the policies whose distributions match the dataset distribution better, as the agent cannot correctly evaluate other policies not covered by the dataset. Many algorithms have been proposed to solve the problem following very different pessimistic learning techniques. Kidambi et al. (2020); Yu et al. (2020, 2021) proposed model-based learning algorithms that learn a world model first and then apply online RL to learn from the world model. Wu et al. (2019) proposes a behavior regularization approach closely related to the regularization considered in the RLHF studies. Fujimoto and Gu (2021); Kostrikov et al. (2021); Kumar et al. (2020); Levine et al. (2020) proposed model-free learning algorithms that don’t need to learn the world model. Cheng et al. (2022) provide theoretical guarantees on the efficiency of their model-free learning methods that the learned policy is always better than the behavior policy used to collect the dataset and can compete with the best policy covered by the dataset. Li et al. (2024) mathematically interprets the essence of pessimistic learning algorithms.
@@ 2.2 Offline Preference-Based Reinforcement Learning
Zhan et al. (2023); Zhu et al. (2023) theoretically investigate the problem of pessimistic learning in the offline PBRL setting. They propose a method that is guaranteed to learn near-optimal policy depending on the dataset, but it remains unknown how to implement these algorithms in practice. Stiennon et al. (2020); Ouyang et al. (2022) studies the problem of language model fine-tuning, where humans provide the preference labels. These works, also known as RLHF, do not apply to the general RL setting. These works also only consider behavior regularization for pessimistic learning, which is shown to be less efficient than other advanced pessimistic learning techniques.
A line of research studies a variant of the offline learning setting. In this setting, the learning agent can access a preference dataset consisting of preference labels over trajectory pairs and a demonstration dataset consisting of only trajectories. Kim et al. (2023) proposes to use a transformer architecture to approximate the reward model. Hejna and Sadigh (2024) proposes to adapt the IQL learning framework to the preference-based learning setting. Zhang et al. (2023) proposes to learn a preference model instead of a reward model and then learn a policy aligned with the preference model.
Another line of research Sadigh et al. (2017); Shin et al. (2021, 2023) studies the PBRL problem in the active learning setting. In this case, the learning starts from a pre-collected behavior dataset with no preference label. During training, the agent can query an expert to provide preference feedback on a pair of trajectories sampled from the behavior dataset by the agent in an online manner. Then, the agent learns a reward model from the preference feedback and applies RL algorithmss to learn from the reward model.
@@ 3.1 Reinforcement Learning
First, we introduce the general RL framework, where an agent interacts with an environment at discrete time steps. The environment is characterized by a Markov Decision Process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{D}\}$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{R}$ is the reward function, and $\mathcal{D}$ is the state transition dynamics. At each time-step, the environment is at a state $s\in\mathcal{S}$, and the agent takes an action $a\in\mathcal{A}$, and then the environment transits to the next state $s^{\prime}\sim\mathcal{D}(\cdot|s,a)$ with an instant reward $r=R(s,a)$. Without loss of generality, we assume the rewards are bounded in $[-1,1]$. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ represents a way to interact with the environment by sampling an action from the distribution given by $\pi(s)$ at a state $s$. Given an initial state $s_{0}$, the performance of a policy is evaluated through its discounted cumulative reward $J(\pi)=\sum_{t=0}^{\infty}\mathbb{E}[\gamma^{t}R(s^{t},a^{t})|a^{t}\sim\pi(s^{t})]$, where $\gamma$ is a discount factor. In the standard RL framework, the learner can observe scalar reward feedback for a state action, which is not the case for the preference-based setting considered in our work.
@@ 3.3 Offline PBRL Setting
In this work, we study the problem of offline PBRL. Here, the environment is still characterized by an MDP =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{D}\}$, but the learner cannot directly interact with the environment or observe any reward signals from $\mathcal{R}$. The learner knows the state and action spaces $\mathcal{S},\mathcal{A}$ and can access a pre-collected dataset $D$ of preference feedback. The dataset $D$ contains $N$ tuples of data, and each tuple consists of a pair of trajectories and a preference label. The preference labels in the dataset are generated by a preference model $P$ unknown to the learning agent. For each pair of trajectory $(\tau_{1},\tau_{2})$, the preference model randomly generates a preference signals $\sigma\in\{1,2\}$ through a Bernoulli trial with a probability $p=P(\tau_{1},\tau_{2})$. The preference signal $\sigma$ represents the index of the preferred trajectory. For convenience, we call the preferred trajectory ‘the chosen trajectory’ and the other trajectory ‘the rejected trajectory.’ Wang et al. (2023) considers a different setting where multiple preference labels are generated for each pair of trajectories, and we will extend our main method to this setting in the experiments. Note that the preference model $P$ is determined by a link function $f$ and the reward function $\mathcal{R}$ of the environment, and both should be unknown to the learner in principle. It is common among recent works that assume $f$ to be a sigmoid function, but in this work, our method does not require the knowledge of $f$. At a high level, the agent’s goal is to reliably learn a policy from the preference dataset $D$ with a high cumulative reward based on the underlying reward function $\mathcal{R}$.
@@ 4 Bridging Preference and Reward-Based Offline Reinforcement Learning
To utilize existing efficient reward-based RL algorithms for PBRL, the most popular choice among recent PBRL studies is reward modeling (Ouyang et al., 2022). However, to bridge the gap, learning a complete reward model at the state actions not in the dataset is unnecessary. All that is needed is to generate the reward labels for the state actions in the dataset, and the reward-based offline RL algorithms will do pessimistic learning and inference on the environment model, such as reward, at unseen state actions during training. Therefore, our task is to find the reward labels that contain the closest information to the preference labels.
@@ 4.1 Optimal Reward Labeling
First, we introduce the basic metric to evaluate how well the rewards labels interpret the preference labels. Given an offline preference dataset $D=\{(\tau_{1}^{i},\tau_{2}^{i},\sigma^{i})\},i\in[N]$, let $(s^{i}_{j,t},a^{i}_{j,t}),j\in[2],t\in[T]$ be the $t^{th}$ state-action pairs of trajectory $j$ from the $i^{th}$ data tuple, and let $r^{i}_{j,t}$ be the reward label for this state action pair. For convenience, we denote $\bar{\sigma}^{i}$ as the index for the rejected trajectory. Let $f$ be the link function between the rewards and the preference, the probability of preferring the chosen trajectory $\tau_{\sigma^{i}}^{i}$ over $\tau_{\bar{\sigma}^{i}}^{i}$ predicted by the rewards label is $p=f(\sum_{t}r^{i}_{\sigma^{i},t}-\sum_{t}r^{i}_{\bar{\sigma}^{i},t})$. Then, a monotonically decreasing loss function $L:[0,1]\rightarrow\mathbb{R}$ can be defined based on the probability that the rewards predict the chosen trajectory. One can use the total prediction loss on preference to evaluate the quality of the reward labels. We denote $F(\cdot)=L(f(\cdot))$ as the link-loss function to represent the combination. Note that the loss and link functions are monotonically decreasing and increasing, respectively. So, the link-loss function is monotonically decreasing. This link-loss function has been widely considered in recent PBRL studies that consider reward modeling. For reward modeling, the reward labels in the loss function are replaced by the reward predictions by the reward model at the corresponding state actions. In these works, the common choice for $f$ is the sigmoid function, and $L$ is the negative log of the KL divergence.
The reward labels with minimal prediction loss are the ones that have information closest to the preference dataset among all possible reward labels. Therefore, our method uses optimal reward labels that induce the minimal prediction loss on the preference dataset $r^{*}=\arg\min_{r}\sum_{i\in[N]}F(\sum_{t}r^{i}_{\sigma^{i},t}-\sum_{t}r^{i}_{\bar{\sigma}^{i},t}))$. In this way, we can transform preference labels into scalar reward labels with minimal information loss. In this work, we consider the most popular case where one preference label is given for each pair of trajectories. In this case, it is straightforward to prove that the optimal reward labels are simply the binary reward labels. We empirically study the case of multiple preference labels for the same trajectory pair in the experiments.
@@
Consider a preference dataset with a single preference label for each pair of trajectories. The optimal reward labels for a loss-link function $F$ and a preference dataset $D$ is the solution to the following optimization problem:
$\arg\min_{r}\sum_{i\in[N]}F(\sum_{t\in[T]}r^{i}_{\sigma^{i},t}-\sum_{t\in[T]}r^{i}_{\bar{\sigma}^{i},t})).$
For any valid loss-link function $F$, the solution (the optimal reward labels) to the problem is:
$r^{i}_{j,t}=\begin{cases}+1,&\forall t\in[T],\forall i\in[N],\text{if }j=\sigma^{i}\\ -1,&\forall t\in[T],\forall i\in[N],\text{if }j=\bar{\sigma}^{i}\end{cases}$
The proof for Lemma 4.1 is in the Appendix. Lemma 4.1 says that for all state actions from the chosen trajectories, the optimal reward labels are the maximal rewards; for all state actions from the rejected trajectories, the optimal reward labels are the minimal rewards. Formally, Alg 4.1 shows how to apply an arbitrary offline reward-based RL algorithm to our framework.
Note that we do not require the reward labels to come from a reward model here. The requirement makes no difference when there is no overlap between the trajectories in the dataset. This can be common in practice. For example, in RLHF, each prompt usually samples two different answers to label (Touvron et al., 2023); in continuous control, the state and action spaces are continuous, making it unlikely to visit the same state action twice. Next, we discuss the special case where the same state action appears in different chosen and rejected trajectories, and the reward labels are required to be the same for the same state action from different trajectories. In this case, we can still apply the binary labeling technique from Lemma 4.1, and the labels for the repeated state actions are simply the mean of the binary rewards. However, this is only a sub-optimal solution. In this paper, to incorporate a more general scenario, we mainly consider the dataset whose trajectories have no overlap. In the Appendix, we empirically study the case where the trajectories overlap with each other. The results of the experiments show that the ORL framework has high and similar learning performance in both cases.
{algorithm}
[…]

### v3

# Binary Reward Labeling: Bridging Offline Preference and Reward-Based Reinforcement Learning

@@ Abstract
Offline reinforcement learning has become one of the most practical RL settings. However, most existing works on offline RL focus on the standard setting with scalar reward feedback. It remains unknown how to universally transfer the existing rich understanding of offline RL from the reward-based to the preference-based setting. In this work, we propose a general framework to bridge this gap. Our key insight is transforming preference feedback to scalar rewards via binary reward labeling (BRL), and then any reward-based offline RL algorithms can be applied to the dataset with the reward labels. The information loss during the feedback signal transition is minimized with binary reward labeling in the practical learning scenarios. We theoretically show the connection between several recent PBRL techniques and our framework combined with specific offline RL algorithms. By combining reward labeling with different algorithms, our framework can lead to new and potentially more efficient offline PBRL algorithms. We empirically test our framework on preference datasets based on the standard D4RL benchmark. When combined with a variety of efficient reward-based offline RL algorithms, the learning result achieved under our framework is comparable to training the same algorithm on the dataset with actual rewards in many cases and better than the recent PBRL baselines in most cases.
@@ 1 Introduction
Reinforcement learning (RL) is an important learning paradigm for solving sequential decision-making problems (Sutton & Barto, 2018). Compared to the standard (reward-based) RL that requires access to reward feedback (Schulman et al., 2017), preference-based RL (PBRL) (Wirth et al., 2017) only requires preference feedback over a pair of trajectories, making it more accessible in practice. In the offline learning setting, the agent only needs a pre-collected dataset of preference labels before training, making it even more convenient (Zhu et al., 2023). When humans provide preference labels, PBRL is known as reinforcement learning from human feedback (RLHF) (Stiennon et al., 2020), a popular framework for aligning large language models.
Despite the success in offline PBRL (Kim et al., 2023; Dai et al., 2023; Hejna & Sadigh, 2024), the topic is less studied compared to the offline RL in the standard reward feedback setting (Li et al., 2024; Levine et al., 2020; Fujimoto & Gu, 2021; Kidambi et al., 2020; Wu et al., 2019; Cheng et al., 2022; Tarasov et al., 2024). The critical feature in offline learning is ‘pessimism’ (Li et al., 2024). The learning agent should be pessimistic about the policies whose behaviors are not included in the dataset. Many pessimistic learning algorithms with theoretical insights have been developed for the standard offline RL setting. However, most empirical works for offline PBRL only adopt limited specific pessimistic learning approaches, such as applying certain regularization to the learning policy (Rafailov et al., 2024; Stiennon et al., 2020), which are shown to be less efficient compared to recent SOTA methods in the standard RL setting (Cheng et al., 2022; Tarasov et al., 2024).
Figure 1: Illustration of the binary reward labeling (BRL) framework. Given a dataset consisting of preference feedback as input, the BRL framework labels the dataset with the rewards that best explain the preference signals from the dataset. With a dataset of reward labels, the learning agent can train on the dataset with any efficient offline RL algorithms such as IQL (Kostrikov et al., 2021) or CQL (Kumar et al., 2020) and enjoy high learning efficiency.
@@
  • Standard offline RL algorithms are very different from each other. Model-based algorithms (Kidambi et al., 2020; Yu et al., 2020) learn an environment model and then run online learning algorithms on the model. Model-free algorithms (Cheng et al., 2022; Fujimoto & Gu, 2021) learn a pair of actor-critic models directly to infer the optimal policy without modeling the environment. The framework needs to apply to all kinds of learning algorithms.
  • Preference feedback contains arguably less information than reward feedback. It is impractical to recover the actual rewards from limited preference feedback. The framework needs to make the standard offline RL algorithms work with incomplete information from preference feedback.
In this work, we build a novel framework called ‘BRL’ based on a reward labeling technique that labels the preference dataset with scalar rewards. The critical insight is that we only need to consider transforming feedback signals from preference to reward format while minimizing information loss. The standard offline RL algorithm can handle the pessimistic learning afterward. Note that reward labeling fundamentally differs from reward modeling, which is widely considered in recent PBRL studies (Kim et al., 2023; Stiennon et al., 2020). Reward modeling aims to infer rewards at state actions that are not in the dataset. In contrast, reward labeling focuses on interpreting the rewards for the state actions inside the preference dataset. To minimize the information loss in the transformation, we want to find the optimal reward labels that best explain the state actions in the dataset. We show that a simple binary labeling technique already gives the optimal reward labels in a practical case where the trajectories in the dataset have no overlap. We further empirically show that in the general case where no matter whether there is an overlap between trajectories or not, the learning algorithms achieve better performance on the reward labels given by the binary labeling technique than on those given by the reward modeling technique. In Fig 1, we illustrate how our framework is used. In summary, our contributions are as follows.
  • We propose a general framework to bridge the gap between offline PBRL and standard offline RL. Our framework involves labeling the dataset with the reward that maintains most information in the preference signals. Afterward, one can train on the reward-labeled dataset with any standard offline RL algorithms. Our framework can be easily implemented without modifying the original RL algorithms. One can possibly construct more efficient PBRL algorithms by applying SOTA standard RL algorithms to our framework.
@@ 2.1 Offline Reward-Based Reinforcement Learning
Offline reward-based reinforcement learning, or standard offline RL, is the most popular offline RL setting. The fundamental challenge for offline RL is known as ‘distribution mismatch’ (Levine et al., 2020). It refers to the phenomenon that the data distribution in the dataset may not match the distribution induced by the optimal policy. Due to distribution mismatch, an efficient learner must be ‘pessimistic’ (Li et al., 2024). In general, it says that an agent should rely more on the policies whose distributions match the dataset distribution better, as the agent cannot correctly evaluate other policies not covered by the dataset. Many algorithms have been proposed to solve the problem following very different pessimistic learning techniques. Kidambi et al. (2020); Yu et al. (2020; 2021) proposed model-based learning algorithms that learn a world model first and then apply online RL to learn from the world model. Wu et al. (2019) proposes a behavior regularization approach closely related to the regularization considered in the RLHF studies. Fujimoto & Gu (2021); Kostrikov et al. (2021); Kumar et al. (2020); Levine et al. (2020) proposed model-free learning algorithms that don’t need to learn the world model. Cheng et al. (2022) provide theoretical guarantees on the efficiency of their model-free learning methods that the learned policy is always better than the behavior policy used to collect the dataset and can compete with the best policy covered by the dataset. Li et al. (2024) mathematically interprets the essence of pessimistic learning algorithms.
@@ 2.2 Offline Preference-Based Reinforcement Learning
Zhan et al. (2023); Zhu et al. (2023) theoretically investigate the problem of pessimistic learning in the standard offline PBRL setting, the same as the one considered in this work. They propose a method that is guaranteed to learn near-optimal policy depending on the dataset, but it remains unknown how to implement these algorithms in practice. Stiennon et al. (2020); Ouyang et al. (2022) studies the problem of language model fine-tuning, where humans provide the preference labels. These works, also known as RLHF, do not apply to the general RL setting. These works also only consider behavior regularization for pessimistic learning, which is shown to be less efficient than other advanced pessimistic learning techniques.
A line of research studies a variant of the offline learning setting. Here, the learning agent can access a preference dataset consisting of preference labels over trajectory pairs and a demonstration dataset consisting of only trajectories. Kim et al. (2023) proposes to use a transformer architecture to approximate the reward model. Hejna & Sadigh (2024) proposes to adapt the IQL learning framework to the preference-based learning setting. Zhang et al. (2023) proposes to learn a preference model instead of a reward model and then learn a policy aligned with the preference model.
Another line of research Sadigh et al. (2017); Shin et al. (2021; 2023) studies the PBRL problem in the active learning setting. In this case, the learning starts from a pre-collected behavior dataset with no preference label. During training, the agent can query an expert to provide preference feedback on a pair of trajectories sampled from the behavior dataset by the agent in an online manner. Then, the agent learns a reward model from the preference feedback and applies RL algorithms to learn from the reward model.
@@ 3.1 Reinforcement Learning
First, we introduce the general RL framework, where an agent interacts with an environment at discrete time steps. The environment is characterized by a Markov Decision Process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{D}\}$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{R}$ is the reward function, and $\mathcal{D}$ is the state transition dynamics. At each time-step, the environment is at a state $s\in\mathcal{S}$, and the agent takes an action $a\in\mathcal{A}$, and then the environment transits to the next state $s^{\prime}\sim\mathcal{D}(\cdot|s,a)$ with an instant reward $r=R(s,a)$. Without loss of generality, we assume the rewards are bounded in $[-1,1]$. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ represents a way to interact with the environment by sampling an action from the distribution given by $\pi(s)$ at a state $s$. Given an initial state $s_{0}$, the performance of a policy is evaluated through its discounted cumulative reward $J(\pi)=\sum_{t=0}^{\infty}\mathbb{E}[\gamma^{t}R(s^{t},a^{t})|a^{t}\sim\pi(s^{t})]$, where $\gamma$ is a discount factor. In the standard RL framework, the learner can observe scalar reward feedback for a state action, which is not the case for the preference-based setting.
@@ 3.3 Offline PBRL Setting
This work studies PBRL in the standard offline setting Zhan et al. (2023). Here, the environment is still characterized by an MDP =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{D}\}$, but the learner cannot directly interact with the environment or observe any reward signals from $\mathcal{R}$. The learner knows the state and action spaces $\mathcal{S},\mathcal{A}$ and can access a pre-collected dataset $D$ of preference feedback. The dataset $D$ contains $N$ tuples of data, and each tuple consists of a pair of trajectories and a preference label. The preference labels in the dataset are generated by a preference model $P$ unknown to the learning agent. For each pair of trajectory $(\tau_{1},\tau_{2})$, the preference model randomly generates a preference signals $\sigma\in\{1,2\}$ through a Bernoulli trial with a probability $p=P(\tau_{1},\tau_{2})$. The preference signal $\sigma$ represents the index of the preferred trajectory. For convenience, we call the preferred trajectory ‘the chosen trajectory’ and the other trajectory ‘the rejected trajectory.’ Wang et al. (2023) considers a different setting where multiple preference labels are generated for each pair of trajectories, and we will extend our main method to this setting in the experiments. Note that the preference model $P$ is determined by a link function $f$ and the reward function $\mathcal{R}$ of the environment, and both should be unknown to the learner in principle. It is common among recent works that assume $f$ to be a sigmoid function, but in this work, our method does not require the knowledge of $f$. At a high level, the agent’s goal is to reliably learn a policy from the preference dataset $D$ with a high cumulative reward based on the underlying reward function $\mathcal{R}$.
@@ 4 Bridging Preference and Reward-Based Offline Reinforcement Learning
To utilize existing efficient reward-based RL algorithms for offline PBRL, we consider an information translation approach that is applicable to universal reward-based RL algorithms. By assigning a reward label for each state-action in the dataset, any reward-based RL algorithm can be trained on the new dataset with the reward labels. Therefore, our task is to find the reward labels that contain the closest information to the preference labels. Note that it is a common approach to train a reward model (also known as reward modeling) and then use the reward model to generate the reward labels Christiano et al. (2017), but this is not necessary in our case as we only need to assign reward labels to the state-actions in the dataset.
@@ 4.1 Optimal Reward Labeling
First, we introduce the basic metric to evaluate how well the rewards labels interpret the preference labels. Given an offline preference dataset $\mathcal{D}=\{(\tau_{1}^{i},\tau_{2}^{i},\sigma^{i})\},i\in[N]$. Without loss of generality, we assume $\sigma^{i}\equiv 1$. That is, the first trajectory in each pair is always the chosen trajectory, and the second ones are always the rejected trajectories. In this case, we simplify the notation of the dataset as $\mathcal{D}=\{\tau^{i}_{1}\succ\tau^{i}_{2}\},i\in[N]$. Let $(s^{i}_{j,t},a^{i}_{j,t}),j\in[2],t\in[T]$ be the $t^{th}$ state-action pairs of trajectory $j$ from the $i^{th}$ data tuple, and let $r^{i}_{j,t}$ be the reward label for this state action pair. In addition, the reward labels should be generated using the same reward model. In other words, there exists a reward function $\hat{\mathcal{R}}$, such that $r^{i}_{j,t}=\hat{\mathcal{R}}(s^{i}_{j,t},a^{i}_{j,t})$. Let $f:\mathbb{R}\rightarrow[0,1]$ be the monotonically increasing link function between the rewards and the preference, the probability of preferring the chosen trajectory $\tau_{1}^{i}$ over $\tau_{2}^{i}$ predicted by the rewards label is $p=f(\sum_{t}r^{i}_{1,t}-\sum_{t}r^{i}_{2,t})$. Then, a monotonically decreasing loss function $L:[0,1]\rightarrow\mathbb{R}$ can be defined based on the probability that the rewards predict the chosen trajectory. One can use the total prediction loss on preference to evaluate the quality of the reward labels. We denote $F(\cdot)=L(f(\cdot))$ as the link-loss function to represent the combination. Note that the loss and link functions are monotonically decreasing and increasing, respectively. So, the link-loss function is monotonically decreasing. This link-loss function has been widely considered in recent PBRL studies that consider reward modeling. For reward modeling, the reward labels in the loss function are replaced by the reward predictions by the reward model at the corresponding state actions. In these works, the common choice for $f$ is the sigmoid function, and $L$ is the negative log of the KL divergence Christiano et al. (2017) .
The reward labels with minimal prediction loss have information closest to the preference dataset among all possible reward labels. Formally, the optimal reward labels are defined as follows.
Definition 4.1 (Optimal reward label).
Given a preference dataset $D=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$. Let $\mathbf{r}$ be the reward labels for the dataset, and $r^{i}_{j,t}$ be the reward label for the state-action pair of trajectory $j$ at step $t$ of the $i^{\text{th}}$ data tuple. Let $F:\mathbb{R}\rightarrow\mathbb{R}$ be a link-loss function to represent the prediction loss of a reward label on a preference label. The optimal reward labels for this dataset are the solution to the optimization problem as below:
$\arg\min_{r}\sum_{i\in[N]}F(\sum_{t\in[T]}r^{i}_{1,t}-\sum_{t\in[T]}r^{i}_{2,t}))$ (1)
$s.t.\exists\hat{\mathcal{R}}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1],r^{i}_{j,t}=\hat{\mathcal{R}}(s^{i}_{j,t},a^{i}_{j,t}).$
Here $(s^{i}_{j,t},a^{i}_{j,t}),j\in[2],t\in[T]$ are the $t^{th}$ state-action pairs of trajectory $j$ from the $i^{th}$ data tuple
To minimize the information loss during the feedback signal transition, one should use the optimal reward to re-label the dataset. In general cases, finding the exact solution can be complicated, and it is common to use a deep neural network to approximate the reward function Christiano et al. (2017) that minimizes the prediction loss. However, in practice, the same state-action pairs will usually not appear multiple times in the dataset. For example, in RLHF, each prompt usually samples two different answers to label (Touvron et al., 2023); in continuous control, the state and action spaces are continuous, making it unlikely to visit the same state action twice. Note that the reward model constraint requires the reward labels for the same state-action to be the same, and this makes no difference if each state-action is unique in the dataset. In this case, we can directly derive the exact solution to the optimal reward labels as shown in Lemma 4.2 below.
@@
Consider a preference dataset $D=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$. If each state-action pair is unique in the dataset, the optimal reward labels are:
$r^{i}_{j,t}=\begin{cases}+1,&\forall t\in[T],\forall i\in[N],\text{if }j=1\\ -1,&\forall t\in[T],\forall i\in[N],\text{if }j=2\end{cases}$ (2)
The proof for Lemma 4.2 is in the Appendix. Lemma 4.2 says that in the typical cases where there is no overlap between the trajectories, the optimal reward labels are binary signals: for all state actions from the chosen trajectories, the optimal reward labels are the maximal rewards; for all state actions from the rejected trajectories, the optimal reward labels are the minimal rewards. Formally, Alg 1 shows how to apply an arbitrary offline reward-based RL algorithm to our framework.
Next, we discuss the general case where the same state action appears in different chosen and rejected trajectories. As mentioned earlier, a common approach is to approximate the solution through reward modeling with a deep neural network. Alternatively, one can still apply the binary labeling technique from Lemma 4.2, and the labels for the repeated state actions in effect are the mean of the binary rewards. This is only a sub-optimal solution. However, in Section 5, our empirical results show that such a simple binary labeling method is efficient in both overlapped and no overlap cases and is generally more efficient than the reward modeling method. Note that another advantage of BRL compared to reward modeling is that it does not require learning a reward model.
Input: preference dataset, offline reward-based RL algorithm Alg
1. Label the state-actions in the dataset with the optimal reward labels according to Eq 2. The resulting dataset with reward labels is $\mathcal{D}_{R}$.
2. Run algorithm Alg on the reward dataset $\mathcal{D}_{R}$ to learn a policy $\pi$.
Output: $\pi$.
Algorithm 1 Binary Reward Labeling for Offline PBRL